OpenStudy (anonymous):

Solve for y: x + 3y > 9.

OpenStudy (foreverandalways13):

@uri

OpenStudy (foreverandalways13):

@Luigi0210

OpenStudy (anonymous):

Solve for y means to isolate the y by using inverse operations.

OpenStudy (anonymous):

Inverse operations: multiplication <-> division addition <->subtraction

OpenStudy (anonymous):

x + 3y > 9 3y > 9.-x \[y > \frac{9}{3}- \frac{x}{3} \rightarrow y > 3- \frac{x}{3} \]

OpenStudy (anonymous):

ok so i got 3y>9-x what do i do now

OpenStudy (anonymous):

Remember that the 3 is being multiplied by the y. You need to do the inverse operation to separate it. What you do to one side, you must do to the other.

OpenStudy (anonymous):

multiply by 3

OpenStudy (anonymous):

\(3 \times y = 9 - x\) Look at it in expanded form...

OpenStudy (anonymous):

i got it is y>1/3x-3

OpenStudy (anonymous):

Right idea, but you made a mistake, can you find it?

OpenStudy (anonymous):

i put 1/3

OpenStudy (anonymous):

You reversed the 9 and the negative x, which would be ok if you kept the negative sign with the x.

OpenStudy (anonymous):

9-x = -x+9

OpenStudy (anonymous):

ohhh i see i corrected it :) thanks